OpenStudy (anonymous):

what is the d value for the following sequence? 6, 10, 14, 18, 22

OpenStudy (amistre64):

the difference between each number is ....

OpenStudy (anonymous):

thank you

OpenStudy (amistre64):

your welcome, but can you determine how much you have to add to the first term to get to the next term? 6 + d = 10

OpenStudy (anonymous):

here is another, find the four arithmetic means between 12 and -13

OpenStudy (amistre64):

add em up and divide by 4 would be my first idea for that one

OpenStudy (anonymous):

k thanks

OpenStudy (amistre64):

to find the mean we add and divide by 2; to find half of that we divde in 2 again to get a dovode by 4

OpenStudy (amistre64):

divide lol

OpenStudy (anonymous):

oh ok. i was like wait.... what....

OpenStudy (anonymous):

the answer options are not making any sense to me....

OpenStudy (amistre64):

essentially, you want the range of the interval divided up evenly into 4 divisions 12 - 13 = 1 ; 1/2 = 1/2 is the midpoint or average. 1/2 + 12 = 12 1/2 ; divded in half is 6 1/4 soo, we should have them as: 1/2 +- 6 1/4 what are your choices?

OpenStudy (amistre64):

i dropped a negative ... opps

OpenStudy (anonymous):

6, -1, -3, -8 6,2,-3,-8 7,-1,-3,-8 7,2,-3,-8

OpenStudy (amistre64):

well, -3 and -8 are common to all the answers so they can be ignored as far as narrowing down options

OpenStudy (amistre64):

hmm, there seems to be a little something i forgot; to find the 4 arith. means; we add up the ends and divide by 4+1 .... for some reason

OpenStudy (amistre64):

not add, but subtract to find the range -13 - 12 = -25 -25/5 = -5 add 5 starting at the left interval to get to the next means -13 -8 -3 2 7 12

OpenStudy (amistre64):

looks like that matched the last one

OpenStudy (amistre64):

seems like alot of work for something you never ever use :)

OpenStudy (anonymous):

cool, thanks. im in an online class and the give the the quizzes before the information, but the information doesnt help anyways cause they expect me to know how to do other forms of math that i have long since forgotten...

OpenStudy (amistre64):

im not fond of online classes

OpenStudy (anonymous):

as long as its simple math or a history class, im good. took me one day to do online economics semester one

OpenStudy (amistre64):

good luck with it ;)

OpenStudy (anonymous):

thanks. gosh, NONE OF THIS MAKES ANY SENSE! bleh....